OpenStudy (anonymous):

Medal!! In which direction must the graph of f(x) = x2 be shifted to produce the graph of g(x) = (x - 8)2?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

Right down up left

OpenStudy (skullpatrol):

$$\Huge f(x) = x^2$$$$\Huge g(x) = (x - 8)^2$$

OpenStudy (skullpatrol):

Graph them and find out.

OpenStudy (skullpatrol):

Just use x= -1, x= 0 and x = 1 in both functions

OpenStudy (skullpatrol):

see what happens for yourself

OpenStudy (anonymous):

this problem wants to see if you know how the values of your function will change if you subtract a quantity from you x values.

OpenStudy (anonymous):

can you just tell me the direction of it

OpenStudy (skullpatrol):